OpenStudy (anonymous):

What is the inverse of f(x) = 3x + 4?

OpenStudy (nick88888888):

Think of it this way: what does f do? It multiplies by 3 then subtracts 4. So the inverse does the opposite in the reverse order; i.e. adds 4, then divides by 3. So the inverse function is f ˉ¹(x) = (x+4)/3

OpenStudy (scorcher219396):

That's not what I got... The way i learned was to swap your x's and y's in the equation (since the inverse of a graph is one reflected over the identity line, aka swapping the x's and y's) So you'd have x=3y+4, then solve for y -3y=4-x y=(x-4)/3

OpenStudy (nick88888888):

ummm Heres another way to put it, Basically, you are switching the 'x's and the 'y's, and then isolating to solve for x: x= 3y-4 3y= x+4 y= (x+4)/3 Therefore y^-1 (that's how you write inverse) =(x+4)/3

OpenStudy (nick88888888):

I prefer the 1st one i did, its just easier in my opinion, there's always a chance to do it more then 1 way :P

OpenStudy (scorcher219396):

The original there is wrong, it should say x=3y+4 not -4